PHY2054 - Physics 2 -Spring 2019 Hi SamirSign Out Help sics HW 2: Ch 20 47 Problem 20.50 8of 10> Part A Review 1 Constants Periodic Table What is the electrie force on a proton 50 fm from the surtace of the nucleus? Express your answer in newtons The nucieus of a Xe atom (an isotope of the eiement xenon with mass 125 u) is 6.0 fm in diameler. It has 54 protons and chargeq+54 (1 fm 1 temometer 1 x 10m.) Hint: Trest What is the protionsacceleration? 8 9 O. 6
Answer #1

distance between cnetr to proton, d = r + h

= (d/2) + h

= 6/2 + 5

= 8fm

= 8*10^-15 m

a) F = k*Q*q/d^2

= 9*10^9*54*1.6*10^-19*1.6*10^-19/(8*10^-15)^2

= 194.4 N

b) a = F/m

= (194.4/1.67*10^-27)

= 1.16*10^29 m/s^2